OpenStudy (anonymous):

f(x) = (x + 2)2 and g(x) = x2 − (x + 2)2. Find f(x) − g(x).

OpenStudy (anonymous):

You must first factor through both equations and you get f(x) = 4+2x and g(x) = x^2 -4-2x now find the sum